# The two-dimensional Rashba-Holstein model

### v1

@@ Abstract
In this work, we investigate the impact of Rashba spin-orbit coupling (RSOC) on the formation of charge-density wave (CDW) and superconducting (SC) phases in the Holstein model on a half-filled square lattice. Using unbiased finite-temperature Quantum Monte Carlo simulations, we go beyond mean-field approaches to determine the ground state order parameter as a function of RSOC and phonon frequency. Our results reveal that the Rashba metal is unstable due to particle-hole instabilities, favoring the emergence of a CDW phase for any RSOC value. In the limit of a pure Rashba hopping, the model exhibits a distinct behavior with the appearance of four Weyl cones at half-filling, where quantum phase transitions are expected to occur at strong interactions. Indeed, a quantum phase transition, belonging to the Gross-Neveu Ising universality class between a semi-metal and CDW emerges at finite phonon frequency dependent coupling $\lambda_{c}$. In the antiadiabatic limit we observe an enhance symmetry in the IR that unifies SC and CDW orders. These results advance our understanding of competing CDW and SC phases in systems with spin-orbit coupling, providing insights that may help clarify the behavior of related materials.
@@ II.1 The model
where sums run over a two-dimensional square lattice under periodic boundary conditions, with $\langle\mathbf{i},\mathbf{j}\rangle$ denoting nearest neighbors sites. Here, we use the standard second quantization formalism, with $c_{\mathbf{i},\sigma}^{\dagger}$ ($c_{\mathbf{i},\sigma}$) being creation (annihilation) operators of electrons in a given site $\mathbf{i}$ with spin $\sigma(=\uparrow,\downarrow)$, while $n_{\mathbf{i},\sigma}\equiv c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{i},\sigma}$ describe the number operators. The first term on the right hand side of Eq. (1) corresponds to the non spin-flip contribution to the kinetic energy of electrons, with $t$ being the hopping integral, while the second term is the chemical potential $\mu$. The third term describes the spin-orbit contribution, of strength $\alpha$, with $[\hat{\sigma}^{\gamma}]_{\sigma,\sigma^{\prime}}$ defining the elements of Pauli matrices ($\gamma=x$, $y$ and $z$). The fourth term denotes the local harmonic oscillators (HO) of mass $M$ and frequency $\omega_{0}$, with $\hat{X}_{\mathbf{i}}$ and $\hat{P}_{\mathbf{i}}$ being the conjugate displacement and momentum operators. The last term in Eq. (1) describes the on-site electron-phonon coupling of strength $g$. Hereafter, we define the lattice constant and the HO’s masses as unities, while setting the energy scale by the hopping integral, $t$.
@@
Before proceeding, it is worthy discussing the limit cases. For instance, the occurrence of long-range order in the pure Holstein model ($\alpha=0$) has been under intense investigation over the past years, at different lattices and fillings. In particular, in the half-filled square lattice, there are strong evidences that support the emergence of CDW for any EPC $g>0$, at any frequency $\omega_{0}$eber and Hohenadler (2018); Hohenadler and Batrouni (2019); Costa et al. (2020), different from the 1D case, which exhibits a quantum phase transition Jeckelmann et al. (1999); Hohenadler and Fehske (2018). In the antiadiabatic limit, the phonon degree of freedom can be integrated out Berger et al. (1995), and the Holstein model is mapped onto the attractive Hubbard model, with effective dynamic electron-electron interaction being $U_{\rm eff}(\omega)=-\frac{g^{2}/\omega^{2}_{0}}{1-(\omega/\omega_{0})^{2}}$. We define $g^{2}/\omega^{2}_{0}\equiv\lambda$, which is the energy scale for the polaron formation, and sets the effective attractive interaction strength in this limit. Notice that the mapping into the attractive Hubbard model leads to a coexistence between CDW and SC for any $|U|>0$, enforced the $SU(2)$ symmetry.
The most relevant effect of the inclusion of a RSOC is the spin-dependent momentum shift, which splits the spin textures, with well-defined chiralities, as illustrated in Fig. 1 (b). It results in a (chiral) splitting of the band structure, shifting the van Hove singularity away from the half-filling, as depicted in Fig. 1 (c). As discussed below, such a shift preserves the Fermi surface nesting, creating particle-hole instabilities Kawano and Hotta (2023); Kubo (2024). Figure 1 (c) also provides another interesting issue: the occurrence of four Dirac cones; two at the Fermi level [$(\pi,0)$ and $(0,\pi)$], and two others above and below the Fermi level [at $(\pi,\pi)$ and $(0,0)$, respectively], as illustrated in Fig. 1 (d). The position of the Dirac cones may change depending on the ratio $\alpha/t$, with all cones being at the Fermi level in the limit case of $t/\alpha=0$ (pure Rashba), where a Weyl semimetal (WSM) emerges; see, e.g., Figs. 1 (e) and (f).
As mentioned in the Introduction, there are only a few recent unbiased studies available on two-dimensional strongly interacting systems in the presence of SOC. Reference Tang et al., 2014 showed that the critical temperatures of the attractive Hubbard model is drastically enhanced by the inclusion of a RSOC, while Refs. Rosenberg et al., 2017; Song et al., 2024 investigate how the symmetry of Cooper pair is affected by the RSOC. Both references analyzed the antiadiabatic limit (i.e. for the attractive Hubbard model), and away from the WSM case.
@@ II.2 Symmetries
$P^{-1}\alpha c_{\mathbf{i},\sigma}^{\dagger}P=\overline{\alpha}e^{i\mathbf{i}\cdot\mathbf{Q}}c_{\mathbf{i},\sigma}^{\phantom{\dagger}}$ (2)
@@
$\hat{X}_{\mathbf{i}}\rightarrow-\hat{X}_{\mathbf{i}},\,\,\hat{P}_{\mathbf{i}}\rightarrow-\hat{P}_{\mathbf{i}}.$ (3)
Here $\mathbf{Q}=(\pi,\pi)/a$ corresponds to the antiferromagnetic wave vector. As a consequence, $\mu=0$ corresponds to half-band filling, $\langle n_{\mathbf{i}}\rangle=1$. In the absence of interactions, the energy bands read:
@@
with $\epsilon(\mathbf{k})=-2t\left(\cos(\mathbf{k}\cdot\mathbf{a}_{x})+\cos(\mathbf{k}\cdot\mathbf{a}_{y})\right)$, and $\mathbf{V}(\mathbf{k})=2\alpha\left(-\sin(\mathbf{k}\cdot\mathbf{a}_{y}),\sin(\mathbf{k}\cdot\mathbf{a}_{x}),0\right)$. For the antiferromagnetic wave $\mathbf{Q}$ and at half-band filling, $\mu=0$, the nesting condition is satisfied:
@@
Thereby, in addition to the Cooper instability in e.g. the singlet s-wave channel we observe an antiferromagnetic instability in the particle-hole channel. Owing to the attractive retarded nature of the interaction one will expect charge density wave order to dominate Noack et al. (1991).
The $t=0$ point is special in the sense that the dispersion relation in the vicinity of the chemical potential becomes that of Dirac fermions.
$E(\mathbf{k})=\pm\sqrt{\sin^{2}(\mathbf{k}\cdot\mathbf{a}_{x})+\sin^{2}(\mathbf{k}\cdot\mathbf{a}_{y})}$ (6)
Expanding around the four time reversal momenta $\mathbf{K}=(0,0),(\pi,\pi),(0,\pi)$, and $(\pi,0)$ yields the follwing Dirac Hamiltonain:
@@
$\equiv v_{F}\sum_{\mathbf{p}}c^{\dagger}_{\mathbf{p}}\left(p_{x}\tau^{z}\sigma^{y}-p_{y}\mu^{z}\tau^{z}\sigma^{x}\right)c^{\phantom{\dagger}}_{\mathbf{p}}$
In the above the Pauli matrices $\tau^{\alpha}$, $\mu^{\alpha}$ and $\sigma^{{}^{\alpha}}$ act on the $\tau$, $\mu$ and $\sigma$ indices. While $\sigma$ denotes the spin index, the four eigenstates of $\tau^{z}\mu^{z}$ encode the four Dirac points. Specifically $\mathbf{K}=(0,0)\leftrightarrow\tau=1,\mu=1$, $\mathbf{K}=(0,\pi)\leftrightarrow\tau=1,\mu=-1$, $\mathbf{K}=(\pi,\pi)\leftrightarrow\tau=-1,\mu=1$ and $\mathbf{K}=(\pi,0)\leftrightarrow\tau=-1,\mu=-1$. In the last line we have suppressed the the indices to lighten the notation. $\mathbf{p}$ measures momentum from the Dirac points and the Fermi velocity is given by $2\alpha a$ with $a$ the lattice constant. While on the lattice, spin-orbit symmetry is discrete, it becomes continuous in the continuum limit with generators $(\tau^{x}\sigma^{z},\tau^{y}\sigma^{z},\tau^{z})$ of SU(2).
In the absence of hopping matrix element the Dirac cones are pinned to the chemical potential. In the continuum approximation, the hopping term acts as Dirac cone dependent chemical potential. In particular:
${\cal H}_{0,t}=-4t\sum_{\mathbf{p}}c^{\dagger}_{\mathbf{p}}\tau^{z}\left(\frac{\mu^{z}+1}{2}\right)c^{\phantom{\dagger}}_{\mathbf{p}}.$ (8)
To account for superconducting as well as charge density wave mass terms, we adopt a Bogoliubov basis:
@@
In this Bogoliubov basis, the Rashba Hamiltonian reads:
${\cal H}_{0,\alpha}=v_{F}\sum_{\mathbf{p}}\eta^{\dagger}_{\mathbf{p}}\left(p_{x}\tau^{z}\sigma^{y}\nu^{z}-p_{y}\tau^{z}\mu^{z}\sigma^{x}\right)\eta^{\phantom{\dagger}}_{\mathbf{p}}.$ (10)
As apparent, the s-wave superconducting order parameter $\Delta$ and the charge density order $\Phi$ from a triplett of anti-commuting mass terms:
@@
The three matrices $\Gamma_{i,j}=\frac{i}{2}\left[M_{i},M_{j}\right]$ commute with the Hamiltonian and generate SO(3) rotations between CDW and superconducting orders. Hence provided that the interaction possesses the same symmetry then we expect degenerate superconducting and CDW fluctuations. We will see that this is indeed the case in the anti-adiabatic limit, where the Holstein interaction reduces to an attractive Hubbard-U term.
The Hubbard interaction term enjoys SO(4) symmetry corresponding to SU(2)-spin and SU(2)-$\eta$ transformations that rotate superconducting states onto CDW orders Yang and Zhang (1990). The spin symmetry is clearly broken by the Rashba spin-orbit coupling but at $t=0$ SU(2)-$\eta$ symmetry is present both in the interaction term as well as is the Dirac Hamiltonian. Hence the appropriate Hamiltonian for this state reads:
${\cal H}={\cal H}_{0,\alpha}-U\int_{V}d^{2}\mathbf{x}\sum_{i=1}^{3}\left(c^{\dagger}_{\mathbf{x}}M_{i}c^{\phantom{\dagger}}_{\mathbf{x}}\right)^{2}$ (12)
This Hamiltonain has SO(3) symmetry that render SSC and CDW states degenerate.
At finite phonon frequencies, the instantaneous Hubbard term gives way to a retarded interaction. This breaks the SO(3) symmetry and at half-filling we expect the charge density wave state to be favored Noack et al. (1991).
@@ III Results and discussion
We begin our analysis by considering the case of fixed $\omega_{0}/t=1$ and EPC $\lambda/t=2$, while varying the RSOC parameter, $\alpha$. We recall that the pure Holstein model has been extensively examined for this set of parameters, showing CDW order below a critical temperature $T_{c}=0.174(2)$Batrouni and Scalettar (2019). The leading effects of the Rashba coupling on this charge order are illustrated in Fig. 2, where we show the behavior of the staggered CDW structure factor as a function of the inverse temperature ($\beta=1/T$), for different values of $\alpha$ and system sizes. As the temperature decreases (i.e., as $\beta$ increases), the behavior of $S_{\text{CDW}}(\textbf{Q})$ changes from being weakly dependent on the system size at high temperatures (low $\beta$) to strongly dependent at low temperatures (high $\beta$). This transition occurs approximately at the critical $T_{c}(\alpha)$. However, as we are dealing with finite system sizes, at very low temperatures, the correlation length becomes larger than $L$, resulting in a plateaus in $S_{\text{CDW}}(\textbf{Q})$, as shown in Fig. 2. Therefore, the reduction of the strength of the $S_{\text{CDW}}(\textbf{Q})$ plateuaus as $\alpha$ increases provides a clear evidence that Rashba coupling weakens the CDW phase. Moreover, as discussed in more details below, the energy scale at which $S_{\text{CDW}}(\textbf{Q})$ begins to increase is shifted to larger values of $\beta$, suggesting that the Rashba coupling not only affects the order parameter but also pushes the critical temperature to lower values.
@@
Figure 4: The ground state phase diagram of the Rashba-Holstein model for fixed $\omega_{0}=1$. The solid red line is just a crossover between a robust CDW (magenta region) and a weak CDW (write region) one.
@@
In line with the previous discussion, Fig. 3 (a) demonstrates that, in the thermodynamic limit, $S_{\rm CDW}(\textbf{Q})/L^{2}$ is significantly impacted by the RSOC, leading to nearly vanishing values for $\alpha/t\approx 1.3$ These results allow us to present the CDW order parameter as a function of $\alpha/t$, as shown in Fig. 3 (b). Notice that, for small values of RSOC, the CDW response remains similar to that of the pure Holstein model, but, as $\alpha$ increases, the order parameter is suppressed. At this point, an important remark should be made: although Fig. 3 suggests a zero order parameter at $\alpha\approx 1.3$, we expect an asymptotic approach to zero instead. This expectation is based on the instability of the particle-hole channel from the nesting of the Fermi surface in the noninteracting regime, as discussed in Sec. II.2. Indeed, weak coupling (mean-field) results, presented in Ref. Fontenele et al., 2024, exhibit asymptotic behavior for the CDW order parameter in the RHM. However, such a very low order parameter response is particularly challenging to capture using QMC methods, since it demands very low temperatures.
@@
Figure 5: The CDW correlation ratio $R_{CDW}(\beta)$ as a function of the inverse temperature ($\beta$=1/T), for fixed (a) $\alpha=0.2$ and (b) $\alpha=1.2$, and different lattice sizes. (c) The extrapolation of the crossings of the $R_{CDW}(\beta)$, between $L$ and $L+\Delta L$ system sizes, $T_{c}(L,L+\Delta L)$. Here, we fixed $\Delta L=4$. (d) The extrapolated critical temperatures of the Rashba-Holstein model for fixed $\omega_{0}/t=1$. The red curve is just a guide to the eye.
@@
Then, the ground state of the Rashba-Holstein model is expected to always manifest as a CDW, for any $\alpha\geq 0$. However, the strength of its order parameter (which is proportional to the charge gap at the Fermi level) depends on the RSOC, becoming exponentially small for large values of $\alpha$. Therefore, for practical purposes, it is useful to identify the crossover region where the system goes from a strong CDW regime to a weak CDW (wCDW) one. Here, this crossover is identified as the point where the extrapolation of $S_{\rm CDW}(\textbf{Q})/L^{2}$ vanishes, e.g. for $\lambda/t=2$ and $\omega_{0}=1$, it is $\alpha\approx 1.25\pm 0.05$. Identifying this crossover is important, particularly because the region with such an exponentially small order parameter is unstable to even small perturbations, such as thermal fluctuations or further neighbor hopping terms (that break the nesting of the Fermi surface), which can lead to the occurrence of a plain Rashba metal.
@@
Figure 6: (a) The CDW correlation ratio, $R_{CDW}(\lambda)$, as a function of $\lambda$ in the limit of the pure Rashba hopping, and fixed $\omega_{0}/\alpha=2$, $\beta=L$, and $\alpha=1$. (b) The extrapolation of the crossings of the $R_{CDW}(\beta)$, between $L$ and $L+\Delta L$ system sizes. Here, we fixed $\Delta L=4$. (c) The phase diagram of the Rashba-Holstein model in the limit of a pure Rashba hopping. The coexistence between CDW and SC occurs only at the limit of $\omega_{0}\to\infty$.

### v2

@@ Abstract
In this work, we investigate the impact of Rashba spin-orbit coupling (RSOC) on the formation of charge-density wave (CDW) and superconducting (SC) phases in the Holstein model on a half-filled square lattice. Using unbiased finite-temperature Quantum Monte Carlo simulations, we go beyond mean-field approaches to determine the ground state order parameter as a function of RSOC and phonon frequency. Our results reveal that the Rashba metal is unstable due to particle-hole instabilities, favoring the emergence of a CDW phase for any RSOC value. In the limit of a pure Rashba hopping, the model exhibits a distinct behavior with the appearance of four Weyl cones at half-filling, where quantum phase transitions are expected to occur at strong interactions. Indeed, a quantum phase transition, belonging to the Gross-Neveu Ising universality class between a semi-metal and CDW emerges at finite phonon frequency dependent coupling $\lambda_{c}$. In the antiadiabatic limit we observe an enhanced symmetry in the infrared that unifies SC and CDW orders. These results advance our understanding of competing CDW and SC phases in systems with spin-orbit coupling, providing insights that may help clarify the behavior of related materials.
@@ II.1 The model
where sums run over a two-dimensional square lattice under periodic boundary conditions, with $\langle\mathbf{i},\mathbf{j}\rangle$ denoting nearest neighbors sites. Here, we use the standard second quantization formalism, with $c_{\mathbf{i},\sigma}^{\dagger}$ ($c_{\mathbf{i},\sigma}$) being creation (annihilation) operators of electrons in a given site $\mathbf{i}$ with spin $\sigma(=\uparrow,\downarrow)$, while $n_{\mathbf{i},\sigma}\equiv c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{i},\sigma}$ describe the number operators. The first term on the right hand side of Eq. (1) corresponds to the non spin-flip contribution to the kinetic energy of electrons, with $t$ being the hopping integral, while the second term is the chemical potential $\mu$. The third term describes the spin-orbit contribution, of strength $\alpha$, with $\hat{\sigma}^{\gamma}]_{\sigma,\sigma^{\prime}}$ defining the elements of Pauli matrices ($\gamma=x$, $y$ and $z$). The fourth term denotes the local harmonic oscillators (HO) of mass $M$ and frequency $\omega_{0}$, with $\hat{X}_{\mathbf{i}}$ and $\hat{P}_{\mathbf{i}}$ being the conjugate displacement and momentum operators. The last term in Eq. (1) describes the on-site electron-phonon coupling of strength $g$. Hereafter, we define the lattice constant and the HO’s masses as unities, while setting the energy scale by the hopping integral, $t$.
@@
Before proceeding, it is worthy discussing the limit cases. For instance, the occurrence of long-range order in the pure Holstein model ($\alpha=0$) has been under intense investigation over the past years, at different lattices and fillings. In particular, in the half-filled square lattice, there are strong evidences that support the emergence of CDW for any EPC $g>0$, at any frequency $\omega_{0}$Weber and Hohenadler (2018); Hohenadler and Batrouni (2019); Costa et al. (2020), different from the 1D case, which exhibits a quantum phase transition Jeckelmann et al. (1999); Hohenadler and Fehske (2018). In the antiadiabatic limit, the phonon degree of freedom can be integrated out Berger et al. (1995), and the Holstein model is mapped onto the attractive Hubbard model, with effective dynamic electron-electron interaction being $U_{\rm eff}(\omega)=-\frac{g^{2}/\omega^{2}_{0}}{1-(\omega/\omega_{0})^{2}}$. We define $g^{2}/\omega^{2}_{0}\equiv\lambda$, which is the energy scale for the polaron formation, and sets the effective attractive interaction strength in this limit. Notice that the mapping into the attractive Hubbard model leads to a coexistence between CDW and SC for any $|U|>0$, enforced the $SU(2)$ symmetry.
We also recall that the mean-field analysis of the Holstein model with RSOC, discussed in Ref. Fontenele et al., 2024, (or any electron-phonon model) corresponds to the limit in which the lattice is already distorted. That is, when the kinetic term of the quantum HO’s are neglected. Though interesting, this limit it is essentially classical, as the Hamiltonian does not contain non-commuting terms. Therefore, the mean-field results are only relevant when dealing with low frequencies, but only qualitatively. Treating quantum fluctuations by adding momentum operators is nontrivial,, and the resulting behavior may be unexpected. One could approach it perturbatively using the Wigner-Kirkwood method (or truncated Wigner approximation) Steel et al. (1998); Polkovnikov (2003); Paprotzki et al. (2024), but this requires keeping $\hbar\omega_{0}/k_{\rm B}T$ small. In our work, by contrast, we employ a nonperturbative (unbiased) methodology, which constitutes a significant improvement over any mean-field or perturbative approach. As mentioned in the Introduction, there are only a few recent unbiased studies available on two-dimensional strongly interacting systems in the presence of SOC. Reference Tang et al., 2014 showed that the critical temperatures of the attractive Hubbard model is drastically enhanced by the inclusion of a RSOC, while Refs. Rosenberg et al., 2017; Song et al., 2024 investigate how the symmetry of Cooper pair is affected by the RSOC. Both references analyzed the antiadiabatic limit (i.e. for the attractive Hubbard model), and away from the WSM case.
@@ II.2 Symmetries
We first note that the Rashba term breaks inversion (parity) symmetry. The inversion operation is represented by a unitary operator $\Pi$ with $\Pi\,\mathbf{r}\,\Pi^{-1}=-\mathbf{r}$; in momentum space, $\Pi\,\mathbf{k}\,\Pi^{-1}=-\mathbf{k}$. The spin operators are left unchanged by $\Pi$. Given that the Rashba spin-orbit term in the square lattice is given by
$H_{R}(\mathbf{k})=2\alpha\bigl(\sin k_{y}\,\sigma_{x}-\sin k_{x}\,\sigma_{y}\bigr),$
it is clear that
$\Pi\,H_{R}(\mathbf{k})\,\Pi^{-1}=H_{R}(-\mathbf{k})=-H_{R}(\mathbf{k}),$
showing that $H_{R}$ is odd in $\mathbf{k}$ and not invariant under $\Pi$. Thus, the full Hamiltonian lacks inversion symmetry for any $\alpha\neq 0$.
[…]
@@
$\mathcal{P}^{-1}\,c_{\mathbf{i},\sigma}^{\dagger}\,\mathcal{P}=\eta_{\mathbf{i}}c_{\mathbf{i},\sigma},$ (2)
@@
$\hat{X}_{\mathbf{i}}\rightarrow-\hat{X}_{\mathbf{i}},\quad\hat{P}_{\mathbf{i}}\rightarrow-\hat{P}_{\mathbf{i}},$ (3)
with $\eta_{\mathbf{i}}=e^{i\mathbf{i}\cdot\mathbf{Q}}$, in which $\mathbf{Q}=(\pi,\pi)/a$ corresponds to the antiferromagnetic wavevector. As a consequence, $\mu=0$ corresponds to half-band filling, $\langle n_{\mathbf{i}}\rangle=1$. In the absence of interactions, the energy bands read:
@@
with $\epsilon(\mathbf{k})=-2t\left[\cos(k_{x})+\cos(k_{y})\right]$, and $\mathbf{V}(\mathbf{k})=2\alpha\left[-\sin(k_{y}),\sin(k_{x}),0\right]$. For the antiferromagnetic wave $\mathbf{Q}$ and at half-filling, $\mu=0$, the nesting condition is satisfied as
@@
Thereby, in addition to the Cooper instability in e.g. the singlet s-wave channel, we observe an antiferromagnetic instability in the particle-hole channel. Owing to the attractive retarded nature of the interaction one will expect charge-density-wave order to dominate Noack et al. (1991).
The $t=0$ point is special, in the sense that the dispersion relation in the vicinity of the chemical potential becomes that of Dirac fermions,
$E(\mathbf{k})=\pm 2\alpha\sqrt{\sin^{2}(k_{x})+\sin^{2}(k_{y})}.$ (6)
Expanding around the four time-reversal nodal points $\mathbf{K}=(0,0),(\pi,\pi),(0,\pi)$, and $(\pi,0)$ yields
$\sin(k_{x})\simeq s_{x}(\mathbf{K})\,p_{x},\qquad\sin(k_{y})\simeq s_{y}(\mathbf{K})\,p_{y},$ (7)
with node-dependent signs
$\begin{array}[]{c|cccc}\mathbf{K}&(0,0)&(\pi,\pi)&(0,\pi)&(\pi,0)\\ \hline\cr s_{x}&+1&-1&+1&-1\\ s_{y}&+1&-1&-1&+1\end{array}$
Therefore, it leads to the following Dirac-like Hamiltonian
$\mathcal{H}_{0,\alpha}\simeq v_{F}\sum_{\mathbf{p}}c^{\dagger}_{\mathbf{p}}\Big[s_{x}(\mathbf{K})\,p_{x}\,\sigma^{y}-s_{y}(\mathbf{K})\,p_{y}\,\sigma^{x}\Big]\,c^{\phantom{\dagger}}_{\mathbf{p}},$ (8)
where $\mathbf{p}=(p_{x},p_{y})$ measures momentum around the nodal points and the Fermi velocity is given by $v_{F}=2\alpha$.
We may relabel the nodes as $\mathbf{K}=(0,0)\leftrightarrow\tau=1,\mu=1$; $\mathbf{K}=(0,\pi)\leftrightarrow\tau=1,\mu=-1$; $\mathbf{K}=(\pi,\pi)\leftrightarrow\tau=-1,\mu=1$; and $\mathbf{K}=(\pi,0)\leftrightarrow\tau=-1,\mu=-1$. This allow us to introduce Pauli matrices $\tau^{z}$ and $\mu^{z}$, so that
$s_{x}(\mathbf{K})=\tau^{z}_{\tau,\tau^{\prime}},\qquad s_{y}(\mathbf{K})=\tau^{z}_{\tau,\tau^{\prime}}\mu^{z}_{\mu,\mu^{\prime}},$ (9)
which, in turn, leads to the compact Dirac-like Hamiltonian
@@
$\equiv v_{F}\sum_{\mathbf{p}}c^{\dagger}_{\mathbf{p}}\left(p_{x}\tau^{z}\sigma^{y}-p_{y}\mu^{z}\tau^{z}\sigma^{x}\right)c^{\phantom{\dagger}}_{\mathbf{p}}~.$
In the last line we have suppressed the indices to lighten the notation. While on the lattice, spin-orbit symmetry is discrete, it becomes continuous in the continuum limit with generators $(\tau^{x}\sigma^{z},\tau^{y}\sigma^{z},\tau^{z})$ of $SU(2)$.
Notice that, any finite hopping $t$ shifts the nodal points at $\mathbf{K}=(0,0)$ and $(\pi,\pi)$ away from the half-filling. In the continuum approximation, this hopping term acts as Dirac cone dependent chemical potential. It can be showed by expanding the hopping term around the nodal points, which leads to
${\cal H}_{t,0}=-4t\sum_{\mathbf{p}}c^{\dagger}_{\mathbf{p}}\tau^{z}\left(\frac{\mu^{z}+1}{2}\right)c^{\phantom{\dagger}}_{\mathbf{p}}.$ (11)
To account for superconducting as well as charge-density-wave mass terms, we adopt a Bogoliubov basis
@@
which, leads to
${\cal H}_{0,\alpha}=v_{F}\sum_{\mathbf{p}}\eta^{\dagger}_{\mathbf{p}}\nu^{z}\left(p_{x}\tau^{z}\sigma^{y}-p_{y}\tau^{z}\mu^{z}\sigma^{x}\right)\eta^{\phantom{\dagger}}_{\mathbf{p}}.$ (13)
The $s$-wave SC order parameter $\Delta$ and the CDW one, $\Phi$, form a triplet of anticommuting mass terms
@@
That is, these three masses mutually anticommute, and also anticommutes with the Dirac-like Hamiltonian in Eq. (13), which eventually produces a gapped dispersion. By defining the generators
$\Gamma_{nm}\;\equiv\;\frac{i}{2}\,[M_{n},M_{m}],\qquad n,m=\{1,2,3\},$ (15)
one finds $\Gamma_{nm}=\varepsilon_{nmk}\,M_{k}$, which satisfy the $SO(3)$ algebra and act on the triplet as vector rotations $\Gamma_{ij},M_{k}]=i\,\varepsilon_{ijk}\,M_{i}$. Hence, the pure Rashba term does not favour SC over CDW (or vice-versa), and provided that the interaction possesses the same symmetry then we expect degenerate SC and CDW fluctuations.
At half-filling the attractive Hubbard interaction possesses an $\mathrm{SO}(4)\!\cong\!\mathrm{SU}(2)_{\text{spin}}\!\times\!\mathrm{SU}(2)_{\eta}$ symmetry; the latter rotates the SC state onto the CDW one Yang and Zhang (1990). The spin symmetry is clearly broken by the Rashba spin-orbit coupling, but at $t=0$ SU(2)η symmetry is present both in the interaction term as well as is the Dirac-like Hamiltonian. Hence, as the appropriate Hamiltonian reads
${\cal H}={\cal H}_{0,\alpha}-U\int_{V}d^{2}\mathbf{x}\sum_{i=1}^{3}\left(c^{\dagger}_{\mathbf{x}}M_{i}c^{\phantom{\dagger}}_{\mathbf{x}}\right)^{2}~,$ (16)
then it has SO(3) symmetry that renders SC and CDW states degenerate for any $U$ value. On the other hand, at finite phonon frequencies, the instantaneous Hubbard term gives way to a retarded interaction. This breaks the SO(3) symmetry and may favor CDW order Noack et al. (1991).
@@ III Results and discussion
We begin our analysis by considering the case of fixed $\omega_{0}/t=1$ and EPC $\lambda/t=2$, while varying the RSOC parameter, $\alpha$. We recall that the pure Holstein model has been extensively examined for this set of parameters, showing CDW order below a critical temperature $T_{c}/t=0.174(2)$Batrouni and Scalettar (2019). The leading effects of the Rashba coupling on this charge order are illustrated in Fig. 2, where we show the behavior of the staggered CDW structure factor as a function of the inverse temperature ($\beta=1/T$), for different values of $\alpha$ and system sizes. As the temperature decreases (i.e., as $\beta$ increases), the behavior of $S_{\text{CDW}}(\textbf{Q})$ changes from being weakly dependent on the system size at high temperatures (low $\beta$) to strongly dependent at low temperatures (high $\beta$). This transition occurs approximately at the critical $T_{c}(\alpha)$. However, as we are dealing with finite system sizes, at very low temperatures, the correlation length becomes larger than $L$, resulting in a plateaus in $S_{\text{CDW}}(\textbf{Q})$, as shown in Fig. 2. Therefore, the reduction of the strength of the $S_{\text{CDW}}(\textbf{Q})$ plateuaus as $\alpha$ increases provides a clear evidence that Rashba coupling weakens the CDW phase. Moreover, as discussed in more details below, the energy scale at which $S_{\text{CDW}}(\textbf{Q})$ begins to increase is shifted to larger values of $\beta$, suggesting that the Rashba coupling not only affects the order parameter but also pushes the critical temperature to lower values.
@@
Figure 4: The ground state phase diagram of the Rashba-Holstein model for fixed $\omega_{0}=1$. The solid red line is just a crossover between a robust CDW (magenta region) and a weak CDW (white region) one.
@@
In line with the previous discussion, Fig. 3 (a) demonstrates that, in the thermodynamic limit, $S_{\rm CDW}(\textbf{Q})/L^{2}$ is significantly impacted by the RSOC, leading to nearly vanishing values for $\alpha/t\approx 1.3$ . These results allow us to present the CDW order parameter as a function of $\alpha/t$, as shown in Fig. 3 (b). Notice that, for small values of RSOC, the CDW response remains similar to that of the pure Holstein model, but, as $\alpha$ increases, the order parameter is suppressed. At this point, a remark should be made: although Fig. 3 suggests a zero order parameter at $\alpha\approx 1.3$, we expect an asymptotic approach to zero instead, due to the instability of the particle-hole channel from the nesting of the Fermi surface. Indeed, from a mean-field perspective, the CDW order parameter exhibits asymptotic behavior as $\alpha\to 0$ Fontenele et al. (2024), analogous to the mean-field behavior of the antiferromagnetic order parameter in the repulsive Hubbard-Rashba model Kubo (2024); Kawano and Hotta (2023); in both cases, the ordering occurs at the wavevector $\mathbf{Q}=(\pi,\pi)$. However, in contrast to the repulsive Hubbard-Rashba model – where increasing RSOC can stabilize additional magnetic orders –, our QMC data for the Holstein model do not indicate distinct CDW orderings over the parameter ranges explored. That is, while the RSOC strongly affects spin ordering in the Hubbard model (in particular at strong coupling by means of a Dzyaloshinskii-Moriya interaction Cocks et al. (2012); Radić et al. (2012)), its interplay with lattice degrees of freedom in the Holstein model is insufficient to stabilize alternative CDW patterns.
@@
Figure 5: The CDW correlation ratio $R_{CDW}(\beta)$ as a function of the inverse temperature ($\beta$=1/T), for fixed (a) $\alpha=0.2$ and (b) $\alpha=1.2$, and different lattice sizes. (c) The extrapolation of the crossings of the $R_{CDW}(\beta)$, between $L$ and $L-\Delta L$ system sizes, $T_{c}(L,L-\Delta L)$. Here, we fixed $\Delta L=4$. (d) The extrapolated critical temperatures of the Rashba-Holstein model for fixed $\omega_{0}/t=1$. The red curve is just a guide to the eye.
@@
Then, the ground state of the Rashba-Holstein model is expected to always manifest as a CDW, for any $\alpha\geq 0$. However, the strength of its order parameter (which is proportional to the charge gap at the Fermi level) depends on the RSOC, becoming exponentially small for large values of $\alpha$. Here, we recall that the particle–hole and particle-particle channels exhibit similar logarithmic divergences. Therefore, in the regime of large RSOC, where we expect an exponentially small CDW order parameter, a competition between CDW and SC orders may occur, although we are unable to quantify it by QMC. Thus, for practical purposes, it is useful to identify the crossover region where the system goes from a strong CDW regime to a weak CDW (wCDW) one, which can be marked at the point where the extrapolation of $S_{\rm CDW}(\textbf{Q})/L^{2}$ vanishes. Identifying this crossover is important, particularly because the region with such an exponentially small order parameter is unstable to even small perturbations. For instance, thermal fluctuations or further neighbor hopping terms (that break the nesting of the Fermi surface) can lead to the occurrence of a plain Rashba metal; similarly, a small degree of disorder can lead to SC ordering.
@@
Figure 6: (a) The CDW correlation ratio, $R_{CDW}(\lambda)$, as a function of $\lambda$ in the limit of the pure Rashba hopping, and fixed $\omega_{0}/\alpha=2$, $\beta=L$, and $\alpha=1$. (b) The extrapolation of the crossings of the $R_{CDW}(\beta)$, between $L$ and $L+\Delta L$ system sizes. Here, we fixed $\Delta L=4$. (c) The phase diagram of the Rashba-Holstein model in the limit of a pure Rashba hopping. The coexistence between CDW and SC occurs for large frequencies, and become degenerate in the limit of $\omega_{0}\to\infty$.